Different cash flow. Given the following cash inflow at the end of each year, what is the future value of this cash flow at 4%, 10%, and 14% interest rates at the end of year 7? Year 1 Year 2: Year 3 Years 4 through 6: Year 7: $13,000 $22,000 $31,000 $0 $140,000 What is the future value of this cash flow at 4% interest rate at the end of year 7? (Round to the nearest cent.)

What is the interest rate at the end of year 7 at 4%, 10% and 14% ???
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

At 4%:

A=13000*(1.04)^6+22000*(1.04)^5+31000*(1.04)^4+140,000

=$219,481.13(Approx).

At 10%:

A=13000*(1.1)^6+22000*(1.1)^5+31000*(1.1)^4+140,000

=$243,848.61(Approx).

At 14%:

A=13000*(1.14)^6+22000*(1.14)^5+31000*(1.14)^4+140,000

=$263251.53(Approx).